OpenStudy (anonymous):

What are the coordinates of the focus of the conic section shown below? y^2+16y-4x+4=0

OpenStudy (anonymous):

(y+4)^2=4(x-5) This is a parabola with vertex at(5,-4), and axis of symmetry, y=-4 4p=4 p=1 ans: equation of directrix, x=4

OpenStudy (anonymous):

focus= (6,-4)